OpenStudy (anonymous):

Involve exponential growth or decay. If you have 50 grams of carbon-14 today, how much will be left in 100 years? Note that the half-life of carbon-14 is 5730 years (ie. y(5730) = 25).

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

y = A*(1/2)^(x/H) A = original amount H = halflife y = 50*(1/2)^(x/5730) y = 50*(1/2)^(100/5730) y = ??

OpenStudy (anonymous):

49.39

jimthompson5910 (jim_thompson5910):

49.39880 which turns into 49.40 if you're rounding to 2 decimal places

OpenStudy (anonymous):

okay